Solve: 2x + 3y + 4z = 20, x + y + 2z = 9, 3x + 2y + z = 10
Answer by mananth(16946)

2 x + 3 y + 4 z = 20 -------------- 1
1 x + 1 y 2 z = 9 -------------- 2
3 x + 2 y + 1 z = 10 -------------- 3

consider equation 1 &2 Eliminate y
Multiply 1 by -1 -5
Multiply 2 by 3 4
we get
-2 x + -3 y + -4 z = -20
3 x + 3 y + 6 z = 27
Add the two
1 x + 0 y + 2 z = 7 ------------- 4
consider equation 2 & 3 Eliminate y
Multiply 2 by -2
Multiply 3 by 1
we get
-2 x + -2 y + -4 z = -18
3 x + 2 y + 1 z = 10
Add the two
1 x + 0 y + -3 z = -8 -------------5 5
Consider (4) & (5) Eliminate x
Multiply 4 by -1
Multiply (5) by 1
we get
-1 x + -2 z = -7
1 x + -3 z = -8
Add the two
0 x + -5 z = -15
/ -5
z = 3
1 x + -3 z= -8
1 x + -9 = -8
x = 1
Plug x & z in eq (1)
2 x + 3 y + 4 z = 20 -------------- 1
2 + 3 y + 12 = 20
3 y = 6
y = 2
x= 1 y= 2 z= 3